The sum of two numbers is 59. If you subtract the smaller number from the larger, the answer is 33. What are the numbers?
Found 2 solutions by Alan3354, myidis:
Answer by Alan3354(69443)

The sum of two numbers is 59. If you subtract the smaller number from the larger, the answer is 33. What are the numbers?
---------
One number is the average of 59 & 33.

Answer by myidis(1)

x+y=59
assume y is the smaller number then
x-y=33 => x=y+33 therefore we have
y+y+33=59 ==>y=(59-33)/2 & X+(59*33)/2=59 ==>x=59-((59-33)/2)